# Human foraging strategies flexibly adapt to resource distribution and time constraints

### v3

@@ 1 Introduction
A formal solution to patch-leaving and other foraging problems is the Marginal Value Theorem of optimal foraging theory (Charnov,, 1976), according to which the most profitable moment to leave a patch (“giving-up time”) is when the reward in the current patch is less than the average reward rate of the environment (Constantino and Daw,, 2015). Neurally, the decision to stay or to leave is influenced by the time-linked reward representations at the level of the dACC (Wittmann et al.,, 2016). Furthermore, a recent PET study has demonstrated the role of dopamine, essential for motivational states that drive human decision-making dynamics (Le Heron et al.,, 2019), in regulating behavioral adaptations to changes in the reward environment by weighing reward benefits against temporal costs (Ianni et al.,, 2023).
Various studies reported that humans and many other animal species behave very close, despite not as good as, the optimal behavior suggested by optimal foraging theory (Pyke et al.,, 1977; Zhang et al.,, 2015; Wolfe,, 2013; Webb et al.,, 2024; Garrett and Daw,, 2020). This suboptimality has many potential causes; among them, patch leaving decisions are influenced by previous performance (Hutchinson et al.,, 2008; Lin and von Helversen,, 2023) as well as individual (van Dooren et al.,, 2021), social (Bidari et al.,, 2022), and aging factors (Lloyd et al.,, 2023).
A series of human studies tested how environmental constraints – the most important being resource distribution – affect individuals’ patch-leaving strategies. Using a multiple-target visual search task, Cain and colleagues (Cain et al.,, 2012) demonstrated that foragers adapt to environmental statistics in a way that is consistent with, although not identical to, the optimal formulation: participants search more when expecting a greater amount of targets and modify their expectations in real time by considering the target distributions across trials. In a different experiment, a virtual patch-foraging task was used to assess the influence of environmental richness on foraging decisions (Constantino and Daw,, 2015). Results have shown that although participant’s behavior was in agreement with the optimal policy there was an “overharvesting” tendency, namely, a tendency to harvest a patch longer than it would be needed for reward maximization; see also (Hutchinson et al.,, 2008). Overharvesting has been explained as a strategy for reducing the cost of behavioral variability (Cash-Padgett and Hayden,, 2020), as a result of optimal learning and adaptability (Harhen and Bornstein,, 2023), or as a phenomenon occurring as a consequence of temporal cognition (Kendall and Wikenheiser,, 2022). Indeed, time definitely matters when making patch-leaving decisions; as the MVT states, longer travel times lead to longer patch residence times, which lowers the forager’s opportunity cost of time as well as the environment’s overall rate of reward (Bustamante et al.,, 2023).
@@
Finally, most human foraging studies used simple paradigms, in which rewards are paired to buttons that are simultaneously present on the screen. It is unclear how well their results generalize to foraging scenarios that resemble more closely animal patch-living problems, in which rewards are presented serially, visibility is partial and which engage sophisticated cognitive processes, such as those required to learn and remember reward locations, orienting and navigating in space, etc. (Mobbs et al.,, 2018; Hayden,, 2018; Yoo et al.,, 2021; Cisek and Pastor-Bernier,, 2014). Crucially, these more challenging and ecologically valid scenarios present various sources of uncertainty about resource locations and distributions—for example, how many reward locations can be reached within limited time, where rewards are spatially located, and how to navigate efficiently to reach them—hence creating multiple occasions for learning. Therefore, they allow for novel questions about how foraging strategies adapt to environmental demands over the experiment, a factor often sidestepped in accounts like optimal foraging theory. For example, they enable researchers to study whether participants improve their foraging strategy and performance over time as they reduce uncertainty about key task dimensions, and whether, at some point, they match (or closely approximate) optimal solutions from foraging theory (Allen et al.,, 2024; Gordon et al.,, 2021; Maselli et al.,, 2023).
@@ 2.1 Participants and design
Eligibility criteria were designed to be inclusive, without imposing restrictive measures to ensure a more diverse sample, aiming to enhance the representativeness and generalizability of the study’s findings. All the procedures were approved by the CNR Ethics committee.
@@ 2.2 Experimental task and procedure
In the experiment, participants controlled a virtual astronaut from a third-person perspective, using arrow keys (up, left, down, right) or the letters WASD to navigate. The virtual environment consisted of a central zone and four adjacent areas, labeled North, South, East, and West, where “treasure boxes” (henceforth, boxes) were located (Figure  1). Each box contained a variable number of coins and participants’ goal was to maximize the number of coins collected during the experiment.
We employed a 2x2 within-subject design to investigate the impact of resource distribution in the environment (“rich environment” vs. “mixed environment”) and time availability (“long battery” vs. “short battery”) on participants’ foraging strategies (Figure 1.a). The experiment comprised a total of 2 blocks, each consisting of 20 trials, for a total of 40 trials. To mitigate potential biases, participants were divided into two groups. One group experienced the 20 “long battery” trials in the initial block followed by the 20 “short battery” trials in the subsequent block, while the other group experienced the reverse order. To familiarize participants with the task, we administered 5 training trials before the starting of the first block. Additionally, 1 inter-block training trial was introduced to make participants accustomed to the variation in battery length between blocks.
Figure 1: Task design and screenshots from the game illustrating different stages of a trial. (a) Graphical representation of the task design (2x2) showing the two manipulated variables: resource distribution in the environment, either characterized by all areas being rich (“rich environment”) or by two rich and two poor areas (“mixed environment”), and time availability, lasting for 50 seconds (“long battery”) or 25 seconds (“short battery”). (b) Trial beginning: the virtual astronaut stands in the middle of the central zone, facing the spaceship. The four distinct areas are visible, the temporal bar at the top left of the screen is fully charged, and the coin counter at the top right is empty. (c) Within areas exploration and foraging (“East” area in shown): the more boxes are found and reached, the temporal bar gets shorter (top left) and the coin counter (top right) gets higher. (d) Trial end: The temporal bar is now empty, and the coin counter displays the total number of coins collected during the trial. The previously accessible areas are not accessible anymore. Only the spaceship can be accessed, enabling players to activate the “reload” function and begin a new trial.
@@
In addition to manipulating the distribution of resources in the environment, the experiment also involved the manipulation of foraging time available to participants. Participants experienced either trials with a prolonged foraging duration of 50 seconds (referred to as the “long battery” condition) or trials with a shorter duration of 25 seconds (referred to as the “short battery” condition) during trials. The available time was accessible by observing the temporal bar in the top left of the screen; it gradually shortened as time elapsed (Figure 1.c,d). Time depletion paused when the astronaut was in the central zone but moved on while in the four surrounding areas. Participants could freely exit any area, returning to the central zone by moving towards any boundary of the area they were foraging within.
@@
Note that this experiment provides a rich arena to study key features of foraging scenarios. First, it required virtually navigating to locate and then open the boxes, which were spread in pseudo-random locations in the four areas. Second, it required making stay-or-leave decisions between continuing foraging in the same area versus changing area. In turn, changing area was associated with both motor costs (as it took almost twice as long to change area as it does to continue foraging in the same one) and cognitive costs (keeping track of which areas have been explored and which haven’t). Third, in mixed environment, it required exploring to identify which areas were rich and which were poor. Furthermore, it required adapting foraging strategies not just to reward distribution but also to time limitations, as signaled by battery length. Finally, and importantly, this experiment required managing multiple sources of uncertainty, some of which can be reduced over time (e.g., how many boxes and coins they can collect on average in the “long” and “short battery” conditions, how much time it takes on average to reach the next box in the same area or after changing area, where are boxes located and how to reach them most efficiently) and some that cannot (e.g., which areas are rich or poor in a specific trial). Therefore, this experiment permits studying learning dynamics during foraging and asking whether participants’ performance increases over time – perhaps approximating an optimal forager – as they reduce uncertainty about the task.
@@ 2.4 Optimality of foraging strategies
Figure 2: Example branch of the stochastic decision tree for a mixed environment. The agent begins in the $S$ (south) area which is a rich area, knowing that the next box will contain $7$ coins. Moreover, the $E$ (east) area has already been visited and the agent knows that it is a poor area and that the next box will contain $4$ coins. Moving actions are color-coded as follows: blue/red/yellow/peach arrows for moving to the south/east/west/north areas respectively, and the green arrow to remain in the current area and collect a box. The area that is currently occupied by the agent is shown in green. The action that moves to the south area is forbidden as the agent is currently in the south area. Moving to the east area and staying in the south area and collecting the next box are both deterministic actions, as the outcome can be perfectly predicted by the agent. The remaining two actions, moving to the west and north area, both lead to a stochastic node. Each node has a $50\%$ probability to reveal that the chosen area is rich and $50\%$ probability to reveal that it is poor.
@@ 3.1 Influence of resource distribution and foraging time on participants’ foraging strategies
We also analyzed the number of boxes collected per trial (Figure 3.c; Table S3). The mixed model analysis indicated that the battery length (Coeff.: 5.432; SE: 0.113; t: 48.045; $P$-value: $<0.001$; 2.5, 97.5 ci: [5.211, 5.654]) and the richness of the environment (Coeff.: 0.244; SE: 0.113; t: 2.159; $P$-value: 0.031; 2.5, 97.5 ci: [0.023, 0.466]) significantly affect the dependent variable. Specifically, participants opened a significantly greater amount of boxes in long battery and rich environment trials.
@@ 3.3 Learning dynamics during the experiment
The red line represents the linear regression fit to the data. Focusing on the total number of coins (Figure 5.a), the coefficient of determination $R^{2}=0.59$ indicates that the $59\%$ of the variance in the total number of coins can be explained by the independent variables and that the observed relationship is statistically significant ($P$-value: $<0.001$). This result suggest that as people become more familiar with the task, they collect more coins. Further analysis on the amount of boxes collected in the first area suggests that the increased performance was at least partially determined by a change in patch-leaving strategy across trials (Figure 5.b). Indeed, the scatter plot shows a significant ($R^{2}=0.29$; $P$-value: $<0.001$) and negative trend across trials indicating that people tended to overharvest less, the more they became familiar with the task. Moreover, as participants became more familiar with the task, they collected a significantly greater number of boxes per trial ($R^{2}=0.55$; $P$-value: $<0.001$) (Figure 5.c) and visited more areas ($R^{2}=0.69$; $P$-value: $<0.001$) (Figure 5.e). Training may account for these outcomes, as demonstrated by the fact that participants’ became faster at moving from one box to another during the course of the experiment ($R^{2}=0.67$; $P$-value: $<0.001$) (Figure 5.f). Additionally, participants’ average time to leave a room after collecting the last treasure box within it decreases over time (R² = 0.19, p = 0.005), suggesting that participants became more efficient at navigating and making stay-or-leave decisions during the experiment (Figure S3). Finally, the average coins gained per box did not show a significant variation across trials ($R^{2}=0.07$; $P$-value: $<0.109$) (Figure 5.d). This is likely due to a combination of collecting more coins, which increases the average, and opening more boxes, which decreases it.
@@
We finally considered that in our experiment, participants were divided into two groups: one group performed the “long battery” block first, followed by the short battery block, while the other group performed the two blocks in the reverse order. Given that the group that experienced the long battery block first had twice the time for learning compared to the other group, we aimed to investigate whether they performed better in the second (short battery) block. For this, we divided participants’ data according to the two groups and we performed non parametric Mann-Whitney U tests on each of our six dependent variables (Figure 6). We found that participants who started the experiment with the long battery block earned more coins (Figure 6.a, $P$-value: $<0.019$), opened more boxes (Figure 6.b, $P$-value: $<0.013$) and moved faster between boxes (Figure 6.f, $P$-value: $<0.025$) in the short battery block, compared to participants who started the experiment with the short battery block. Finally, participants who started the experiment with the short battery block visited significantly more areas in the long battery block (Figure 6.e, $P$-value: $<0.049$).

### v4

@@ 1 Introduction
A formal solution to patch-leaving and other foraging problems is the Marginal Value Theorem of optimal foraging theory (Charnov,, 1976), according to which the most profitable moment to leave a patch (“giving-up time”) is when the reward in the current patch is less than the average reward rate of the environment (Constantino and Daw,, 2015). Various studies reported that humans and many other animal species behave very close, despite not as good as, the optimal behavior suggested by optimal foraging theory (Pyke et al.,, 1977; Zhang et al.,, 2015; Wolfe,, 2013; Webb et al.,, 2024; Garrett and Daw,, 2020). This suboptimality has many potential causes; among them, patch leaving decisions are influenced by previous performance (Hutchinson et al.,, 2008; Lin and von Helversen,, 2023) as well as individual (van Dooren et al.,, 2021), social (Bidari et al.,, 2022), and aging factors (Lloyd et al.,, 2023).
A series of human studies tested how environmental constraints – the most important being resource distribution – affect individuals’ patch-leaving strategies. Using a multiple-target visual search task, Cain and colleagues (Cain et al.,, 2012) demonstrated that foragers adapt to environmental statistics in a way that is consistent with, although not identical to, the optimal formulation: participants search more when expecting a greater amount of targets and modify their expectations in real time by considering the target distributions across trials. In a different experiment, a virtual patch-foraging task was used to assess the influence of environmental richness on foraging decisions (Constantino and Daw,, 2015). Results have shown that although participant’s behavior was in agreement with the optimal policy there was an “overharvesting” tendency, namely, a tendency to harvest a patch longer than it would be needed for reward maximization; see also (Hutchinson et al.,, 2008). Overharvesting has been explained as a strategy for reducing the cost of behavioral variability (Cash-Padgett and Hayden,, 2020), as a result of optimal learning and adaptability (Harhen and Bornstein,, 2023), as a result of the the inherent asymmetrical cost, in foregone reward, of understaying, especially at longer travel times (Kacelnik,, 1984), or as a phenomenon occurring as a consequence of temporal cognition (Kendall and Wikenheiser,, 2022). Indeed, time definitely matters when making patch-leaving decisions; as the MVT states, longer travel times lead to longer patch residence times, which lowers the forager’s opportunity cost of time as well as the environment’s overall rate of reward (Bustamante et al.,, 2023).
@@
Finally, most human foraging studies used simple paradigms, in which rewards are paired to buttons that are simultaneously present on the screen. It is unclear how well their results generalize to foraging scenarios that resemble more closely animal patch-living problems, in which rewards are presented serially, visibility is partial and which engage sophisticated cognitive processes, such as those required to learn and remember reward locations, orienting and navigating in space, etc. (Mobbs et al.,, 2018; Hayden,, 2018; Yoo et al.,, 2021; Cisek and Pastor-Bernier,, 2014). Crucially, these more challenging and ecologically valid scenarios present various sources of uncertainty about resource locations and distributions—for example, how many reward locations can be reached within limited time, where rewards are spatially located, and how to navigate efficiently to reach them—hence creating multiple occasions for learning. Therefore, they allow for novel questions about how foraging strategies adapt to environmental demands over the experiment, a factor often sidestepped in accounts like optimal foraging theory. For example, they enable researchers to study whether participants improve their foraging strategy and performance over time (both across and within trials) as they reduce uncertainty about key task dimensions, and whether, at some point, they match (or closely approximate) optimal solutions from foraging theory (Allen et al.,, 2024; Gordon et al.,, 2021; Maselli et al.,, 2023).
@@ 2.1 Participants and design
Eligibility criteria were designed to be inclusive, without imposing restrictive measures to ensure a more diverse sample, aiming to enhance the representativeness and generalizability of the study’s findings. Most participants (26 out of 34) reported not being regular gamers. All the procedures were approved by the CNR Ethics committee.
@@ 2.2 Experimental task and procedure
In the experiment, created with Unity (Unity Technologies, San Francisco, US), participants controlled a virtual astronaut from a third-person perspective, using arrow keys (up, left, down, right) or the letters WASD to navigate. The virtual environment consisted of a central zone and four adjacent areas, labeled North, South, East, and West, where “treasure boxes” (henceforth, boxes) were located (Figure  1). Each box contained a variable number of coins and participants’ goal was to maximize the number of coins collected during the experiment.
We employed a 2x2 within-subject design to investigate the impact of resource distribution in the environment (“rich environment” vs. “mixed environment”) and time availability (“long battery” vs. “short battery”) on participants’ foraging strategies (Figure 1.a). The experiment comprised a total of 2 blocks, each consisting of 20 trials, for a total of 40 trials. To mitigate potential biases, participants were divided into two groups. One group experienced the 20 “long battery” trials in the initial block followed by the 20 “short battery” trials in the subsequent block, while the other group experienced the reverse order. While battery length was manipulated between blocks, the environmental richness was varied within blocks, with each trial being randomly assigned to either a “rich ” or a “mixed ” environment condition. This implies that participants had to check the sensor at the beginning of each trial, to know whether the environment was rich or mixed (but remind that the sensor did not tell participants which areas was rich or poor). To familiarize participants with the task, we administered 5 training trials before the starting of the first block. Additionally, 1 inter-block training trial was introduced to make participants accustomed to the variation in battery length between blocks.
Figure 1: Task design and screenshots from the game illustrating different stages of a trial. (a) Graphical representation of the task design (2x2) showing the two manipulated variables: resource distribution in the environment, either characterized by all areas being rich (“rich environment”) or by two rich and two poor areas (“mixed environment”), and time availability, lasting for 50 seconds (“long battery”) or 25 seconds (“short battery”). (b) Trial beginning: the virtual astronaut stands in the middle of the central zone, facing the spaceship. The four distinct areas are visible, the temporal bar at the top left of the screen is fully charged, and the coin counter at the top right is empty. Note that only the leftmost (bronze-colored) coins are used in this experiment, the other two types of coins are not. (c) Within areas exploration and foraging (“East” area in shown): the temporal bar decreased with time, reflecting the time spent by the astronaut in the area (the temporal bar did not decrease in the central zone). As shown in the top center of the screen, up to five boxes can be opened per area to collect coins. (d) Trial end: The temporal bar is now empty, and the coin counter displays the total number of coins collected during the trial. The previously accessible areas are not accessible anymore. Only the spaceship can be accessed, enabling players to activate the “reload” function and begin a new trial.
@@
In addition to manipulating the distribution of resources in the environment, the experiment also involved the manipulation of foraging time available to participants. Participants experienced either trials with a prolonged foraging duration of 50 seconds (referred to as the “long battery” condition) or trials with a shorter duration of 25 seconds (referred to as the “short battery” condition) during trials. Note that the time required to travel between boxes within the same area, or to travel from one area to another, was kept consistent across all experimental conditions (rich/poor areas, long/short battery conditions). The available time was accessible by observing the temporal bar in the top left of the screen; it gradually shortened as time elapsed (Figure 1.c,d). Time depletion paused when the astronaut was in the central zone but moved on while in the four surrounding areas. Participants could freely exit any area, returning to the central zone by moving towards any boundary of the area they were foraging within. The time cost in this experiment was implicit in the movement required for harvesting boxes and traveling between areas. There was no explicit time delay or cost added for these actions. Participants could revisit any area within a trial; however, boxes that had already been collected did not refill and remained empty.
@@
Note that this experiment provides a rich arena to study key features of foraging scenarios. First, it required virtually navigating to locate and then open the boxes, which were spread in pseudo-random locations in the four areas. Each area contained ten predefined treasure box locations, which remained the same for all the study. However, for each trial, only five (randomly selected) out of these ten locations contained boxes. Note that once all five boxes have been opened in an area, that area becomes empty. Second, it required making stay-or-leave decisions between continuing foraging in the same area versus changing area. In turn, changing area was associated with both motor costs, as it took almost twice as long to change area and collect a box in the new area (average time: 6.10 sec) compared to continuing foraging in the same one (average time: 3.41 sec), and cognitive costs, i.e. keeping track of which areas have been explored and which haven’t. Third, in mixed environment, it required exploring to identify which areas were rich and which were poor. Furthermore, it required adapting foraging strategies not just to reward distribution but also to time limitations, as signaled by battery length. Finally, and importantly, this experiment required managing multiple sources of uncertainty, some of which can be reduced over time (e.g., how many boxes and coins they can collect on average in the “long” and “short battery” conditions, how much time it takes on average to reach the next box in the same area or after changing area, where are boxes located and how to reach them most efficiently) and some that cannot (e.g., which areas are rich or poor in a specific trial). Therefore, this experiment permits studying learning dynamics during foraging and asking whether participants’ performance increases over time – perhaps approximating an optimal forager – as they reduce uncertainty about the task.
@@ 2.4 Optimality of foraging strategies
Figure 2: Example branch of the stochastic decision tree for a mixed environment. The agent begins in the $S$ (south) area which is a rich area, knowing that the next box will contain $7$ coins. Moreover, the $E$ (east) area has already been visited and the agent knows that it is a poor area and that the next box will contain $4$ coins. Moving actions are color-coded as follows: blue/red/yellow/peach arrows for moving to the south/east/west/north areas respectively, and the green arrow to remain in the current area and collect a box. The area that is currently occupied by the agent is shown in green. The bottom left cell corresponds to the South (S) area, the bottom right to East (E), the top left to West (W), and the top right to North (N). The action that moves to the south area is forbidden as the agent is currently in the south area. Moving to the east area and staying in the south area and collecting the next box are both deterministic actions, as the outcome can be perfectly predicted by the agent. The remaining two actions, moving to the west and north area, both lead to a stochastic node. Each node has a $50\%$ probability to reveal that the chosen area is rich and $50\%$ probability to reveal that it is poor. The total time associated with each node reflects the cumulative time spent: initially, one box was collected in the East area ($T_{1}$), followed by a move to the South area and collection of one box there ($T_{2}$), summing to $T_{1}+T_{2}$; each subsequent action adds either $T_{1}$ (staying and collecting another box) or $T_{2}$ (moving to a new area), yielding updated totals such as $2T_{1}+T_{2}$ or $T_{1}+2T_{2}$.
@@ 3.1 Influence of resource distribution and foraging time on participants’ foraging strategies
We also analyzed the number of boxes collected per trial (Figure 3.c; Table S3). The mixed model analysis indicated that the battery length (Coeff.: 5.432; SE: 0.113; t: 48.045; $P$-value: $<0.001$; 2.5, 97.5 ci: [5.211, 5.654]) and the richness of the environment (Coeff.: 0.244; SE: 0.113; t: 2.159; $P$-value: 0.031; 2.5, 97.5 ci: [0.023, 0.466]) significantly affect the dependent variable. However, while the environment richness effect was significant as a main effect in the linear mixed model, pairwise post-hoc contrasts between rich and mixed environments did not reach significance under either battery condition. As a follow-up analysis, we compared the number of coins collected before leaving rich areas (Figure S3.a) and poor areas (Figure S3.b). We found significant differences between the number of coins collected under the different conditions, ruling out the possibility that participants followed a fixed, threshold-driven strategy to decide when to leave areas.
@@ 3.3 Learning dynamics during the experiment
The red line represents the linear regression fit to the data. Focusing on the total number of coins (Figure 5.a), the coefficient of determination $R^{2}=0.59$ indicates that the $59\%$ of the variance in the total number of coins can be explained by the independent variables and that the observed relationship is statistically significant ($P$-value: $<0.001$). This result suggest that as people become more familiar with the task, they collect more coins. Further analysis on the amount of boxes collected in the first area suggests that the increased performance was at least partially determined by a change in patch-leaving strategy across trials (Figure 5.b). Indeed, the scatter plot shows a significant ($R^{2}=0.29$; $P$-value: $<0.001$) and negative trend across trials indicating that people tended to overharvest less, the more they became familiar with the task. Moreover, as participants became more familiar with the task, they collected a significantly greater number of boxes per trial ($R^{2}=0.55$; $P$-value: $<0.001$) (Figure 5.c) and visited more areas ($R^{2}=0.69$; $P$-value: $<0.001$) (Figure 5.e). Training may account for these outcomes, as demonstrated by the fact that participants’ became faster at moving from one box to another during the course of the experiment ($R^{2}=0.67$; $P$-value: $<0.001$) (Figure 5.f). Additionally, participants’ average exit time (i.e., the time elapsed between collecting the last box in an area and exiting that area) decreases over time (R² = 0.19, p = 0.005), suggesting that participants became more efficient at navigating and making stay-or-leave decisions during the experiment (Figure S4). Finally, the average coins gained per box did not show a significant variation across trials ($R^{2}=0.07$; $P$-value: $<0.109$) (Figure 5.d). This is likely due to a combination of collecting more coins, which increases the average, and opening more boxes, which decreases it.
@@
We also considered that in our experiment, participants were divided into two groups: one group performed the “long battery” block first, followed by the short battery block, while the other group performed the two blocks in the reverse order. Given that the group that experienced the long battery block first had twice the time for learning compared to the other group, we aimed to investigate whether they performed better in the second (short battery) block. For this, we divided participants’ data according to the two groups and we performed non parametric Mann-Whitney U tests on each of our six dependent variables (Figure 6). We found that participants who started the experiment with the long battery block earned more coins (Figure 6.a, $P$-value: $<0.019$), opened more boxes (Figure 6.b, $P$-value: $<0.013$) and moved faster between boxes (Figure 6.f, $P$-value: $<0.025$) in the short battery block, compared to participants who started the experiment with the short battery block. Finally, participants who started the experiment with the short battery block visited significantly more areas in the long battery block (Figure 6.e, $P$-value: $<0.049$).
@@
Finally, we asked whether participants could adapt their behavior not only between trials (as considered in the previous analyses) but also within trials. We reasoned that as participants navigated across areas, they could have deliberated more or less depending on their level of uncertainty about resource distribution. To test this idea, we repeated the analysis of participants’exit time (illustrated in Figure S4), but within trials. In mixed environments where the richness of each area is unknown and must be inferred through experience, we expected longer exit times early in the trial—when participants are still unsure about the overall resource distribution—and shorter times later, when uncertainty has been (at least partially) resolved. Conversely, in rich environments where all areas are known to be rich from the start, exit time should not depend on the area visit order, because no additional information is being gathered across areas.
To test this hypothesis, we ran two separate linear mixed-effects models for rich and mixed environments, with exit time as the dependent variable and area visit order as a fixed-effect regressor. The models included random intercepts per subject. The analysis revealed no significant effect of area visit order in the rich environment ($p=0.343$), suggesting that exit time remained stable across areas. In contrast, we found a significant negative effect in the mixed environment ($p=0.01$), indicating that participants exited later areas more quickly than earlier ones (Table 1). These findings support the idea that exit time reflects participants’ evolving certainty about the environment within each trial.
Rich Environment Coeff. SE t 2.5 ci 97.5 ci p-value
Intercept 2.675 0.085 31.615 2.509 2.841 $\bm{<0.001}$
Area visit order -0.030 0.032 -0.948 -0.093 0.032 0.343
Mixed Environment
Intercept 2.726 0.067 40.572 2.594 2.858 $\bm{<0.001}$
Area visit order -0.073 0.028 -2.562 -0.128 -0.017 $\bm{0.01}$
Table 1: Results of the linear mixed-effects models assessing whether exit time (in seconds) changes with the area visit order, separately for rich and mixed environments. A significant negative effect of area visit order is observed only in mixed environments, indicating faster exits from later areas as participants reduce uncertainty about resource distribution.